# Optimally-Weighted Maximum Mean Discrepancy Framework for Continual Learning

### v1

@@ 3 Methodology
$F^{\rm feature}_{\theta_{i},k}$ The $k$-th feature layer of $F_{\theta_{i}}$.
$F^{\rm classifier}_{\theta_{i}}$ Linear classifier of $F_{\theta_{i}}$.
${\mathcal{Z}}^{k}$ Feature space of the outputs of $F^{\rm feature}_{\theta_{i},k}$.
${\mathcal{Z}}^{c}$ Feature space of the last feature layer of $F_{\theta_{i}}$.
$\mathcal{F}_{\rm feature}({\theta_{i}},\cdot,k)$ The feature vector extracted by $F^{\rm feature}_{\theta_{i},k}$.
@@
$\mathcal{F}_{r}({\theta_{i}},{\theta_{i+1}},\cdot)$ A regularization term between the previous model $F_{\theta_{i}}$ and the current model $F_{\theta_{i+1}}$.
$F_{d}(\cdot,\cdot)$ Distance measure function.
@@
$w_{k}$ Adaptive weight for regularization term at the $k$-th layer of the network.
${\tilde{w}}_{k}$ The softmax normalized weight corresponding to $w_{k}$.
3.1 Problem definition
In continual learning, a model is unable to access the complete training dataset simultaneously and instead acquires knowledge from a continuously evolving data stream. A common scenario in continual learning is referred to as class-incremental learning, which encompasses a sequence of $N$ tasks $\{T_{1},T_{2},\dots,T_{N}\}$. Each task $T_{i}$ is linked to a labeled training dataset $D^{s}_{i}=\{({\bf x}^{i}_{j},{\bf y}^{i}_{j})\,|\,j=1,\cdots,N^{s}_{i}\}$, where ${\bf x}^{i}_{j}\subset\mathcal{X}$ and ${\bf y}^{i}_{j}\subset\mathcal{Y}$ represent the $j$-th paired sample from the $i$-th task, with $N^{s}_{i}$ indicating the total number of samples in the training dataset $D^{s}_{i}$. Here, $\mathcal{X}$ and $\mathcal{Y}$ denote the respective spaces of data samples and labels. Additionally, let $D^{t}_{i}=\{({\bf x}^{i}_{j},{\bf y}^{i}_{j})\,|\,j=1,\cdots,N^{t}_{i}\}$ represent the testing dataset for the $i$-th task, where $N^{t}_{i}$ signifies the total number of testing samples associated with the $i$-th task.
In a continual learning framework, the goal of the model is to identify the optimal parameter set $\tilde{\Theta}$ that minimizes the training loss across all tasks ${T_{1},\cdots,T_{i}}$ while learning the current task $T_{i}$, as expressed by the following equation :
@@ 3.2 Multi-Level Feature Matching Mechanism
In the realm of continual learning, numerous studies have advocated for the utilization of knowledge distillation [30, 48, 19] methodologies to mitigate the phenomenon of network forgetting. The core principle of knowledge distillation involves maintaining the previous model as a teacher module and synchronizing the outputs between the teacher and the student module, which is represented by the current classifier. This alignment seeks to minimize substantial alterations in critical parameters of the current model (student) during the acquisition of new tasks [5]. Nevertheless, the majority of existing knowledge distillation techniques primarily focus on deriving a regularization term within the prediction space, often neglecting the semantic feature space. In this paper, we present an innovative approach termed the Multi-Level Feature Matching Mechanism (MLFMM), designed to regulate information flow within the feature space, thereby aiming to avert network forgetting.
@@
Let us formally define the function $F_{\theta_{i}}\colon{\mathcal{X}}\to{\mathcal{Y}}$ as a model parameterized by $\theta_{i}$, which comprises $c$ feature layers denoted as $\{F^{\rm feature}_{\theta_{i},1},\cdots,F^{\rm feature}_{\theta_{i},c}\}$, along with a linear classifier represented by $F^{\rm classifier}_{\theta_{i}}\colon{\mathcal{Z}}^{c}\to{\mathcal{Y}}$. Here, the index $i$ signifies the model’s adaptation during the $i$-th task learning phase, and ${\mathcal{Z}}^{c}$ refers to the feature space corresponding to the final feature layer. For a specified input ${\bf x}$, we can derive the feature vector from a particular feature layer through the following process :
@@
$\mathcal{F}_{\rm feature}({\theta_{i}},{\bf x},k)=\begin{cases}F^{\rm feature}_{\theta_{i},1}({\bf x})&k=1\\ F^{\rm feature}_{\theta_{i},k}(F^{\rm feature}_{\theta_{i},1}({\bf x}))&k=2\\ F^{\rm feature}_{\theta_{i},k}(F^{\rm feature}_{\theta_{i},k-1}(\cdots F^{\rm feature}_{\theta_{i},1}({\bf x})))&3\leq k\leq c\,,\end{cases}$ (2)
@@
where each feature layer $F^{\rm feature}_{\theta_{i},k}$ receives the output from the last layer $F^{\rm feature}_{\theta_{i},k-1}$ and produces the feature vector over the space ${\mathcal{Z}}^{k}$. The whole prediction for the model $F_{\theta_{i}}$ can be expressed as :
${\bf\hat{y}}=F^{\rm classifier}_{\theta_{i}}(F^{\rm feature}_{\theta_{i},k}(F^{\rm feature}_{\theta_{i},k-1}(\cdots F^{\rm feature}_{\theta_{i},1}({\bf x}))))\,,$ (3)
@@
where ${\bf\hat{y}}$ is the prediction for the input ${\bf x}$. In this paper, we adopt the cross-entropy loss function to update the parameter ${\theta_{i}}$, defined by :
${\mathcal{L}}_{s}({\bf y},{\bf\hat{y}})=\sum_{t=1}^{\mathbb{C}}\big{\{}{\bf y}[t]\log({\bf\hat{y}}[t])\big{\}}\,,$ (4)
where $\mathbb{C}$ is the total number of classes.
@@
To mitigate the issue of network forgetting, this study implements a memory buffer ${\mathcal{M}}_{i}$ to retain a limited number of historical examples. In particular, we propose utilizing reservoir sampling for the sample selection process, which offers computational efficiency. Nonetheless, the model’s performance on prior tasks diminishes when the memory capacity is constrained. To tackle this challenge, the proposed MLFMM regulates the updates to the model’s representations during the learning of new tasks. Specifically, upon the completion of the current task $T_{i}$, we preserve and freeze the parameters of all feature layers $\{F^{\rm feature}_{\theta_{i},1},\cdots,F^{\rm feature}_{\theta_{i},c}\}$ during the subsequent task learning $T_{i+1}$. A regularization term is incorporated to minimize the discrepancy in representations between the previous model ${F_{\theta_{i}}}$ and the current model ${F_{\theta_{i+1}}}$ at $T_{i+1}$ :
${\mathcal{F}}_{r}({\theta_{i}},{\theta_{i+1}},{\bf x})=\sum_{t=1}^{c}\big{\{}F_{d}(\mathcal{F}_{\rm feature}({\theta_{i}},{\bf x},t),F_{d}(\mathcal{F}_{\rm feature}({\theta_{i+1}},{\bf x},t))\big{\}}\,,$ (5)
@@
where $F_{d}(\cdot,\cdot)$ is a distance measure function and $c$ is the total number of feature layers. In the following section, we introduce a probabilistic distance to implement Eq. (5).
@@ 3.3 The Maximum Mean Discrepancy based Regularization
${\bf Z}^{j}_{i}=\{\mathcal{F}_{\rm feature}({\theta_{i}},{\bf x}_{1},j),\cdots,\mathcal{F}_{\rm feature}({\theta_{i}},{\bf x}_{b},j)\}\,,$ (13)
Similar, we can form the feature vector ${\bf Z}^{j}_{i+1}$ using the current model $F_{\theta_{i}}$ on the data batch $\{{\bf x}_{1},\cdots,{\bf x}_{b}\}$, expressed as :
${\bf Z}^{j}_{i+1}=\{\mathcal{F}_{\rm feature}({\theta_{i+1}},{\bf x}_{1},j),\cdots,\mathcal{F}_{\rm feature}({\theta_{i+1}},{\bf x}_{b},j)\}\,.$ (14)
Let $P_{{\bf Z}^{j}_{i}}$ and $P_{{\bf Z}^{j}_{i+1}}$ denote the distribution of ${\bf Z}^{j}_{i}$ and ${\bf Z}^{j}_{i+1}$, respectively. Based on the MMD criterion and Eq. (5), we can implement the regularization function when seeing the data batch $\{{\bf x}_{1},\cdots,{\bf x}_{b}\}$ at the new task learning ${\mathcal{T}}_{i+1}$ as :
${\mathcal{L}}_{r}=\sum_{t=1}^{c}\big{\{}{\mathcal{L}}^{e}_{{\rm M}}(P_{{\bf Z}^{t}_{i}},P_{{\bf Z}^{t}_{i+1}})\big{\}}\,.$ (15)
@@
where $P_{D^{s}_{i+1}}$ and $P_{{\mathcal{M}}_{i+1}}$ denote the distribution of $D^{s}_{i+1}$ and the memory buffer ${\mathcal{M}}_{i+1}$ at the $(i+1)$-th task learning, ${\bf z}$ is the corresponding pre-softmax responses (logits). Thus, we have $F_{\theta}(x)\triangleq\text{softmax}(H_{\theta}(x))$. This approach helps the model retain past knowledge while learning new tasks efficiently.
Since the proposed approach can be smoothly applied to the existing continual leanring models, we consider applying our appraoch to DER++. The final objective function, including a regularization term $\mathcal{L}_{r}$, is defined as :
@@ 3.4 Adaptive Regularization Optimization Term
To improve the efficacy of the proposed MMD-based regularization method [47, 49], it is essential to automatically allocate distinct weights to each regularization term. This strategy enables the assignment of varying degrees of significance to the features derived from different layers, thereby accurately reflecting their contributions to the overall learning framework. Specifically, we can establish an adaptive weight for each layer to assess its relevance to the entire learning process. To ensure that the adaptive weights across all layers are normalized and suitably balanced, we incorporate the softmax function.
Let ${\bf w}=\{w_{1},\cdots,w_{c}\}$ be a trainable adaptive weight vector, where $w_{j}$ determines the importance of the $j$-th feature layer. To avoid numerical overflow, we propose to normalize all weights using the following equation :
$\tilde{w}_{j}=\frac{e^{w_{j}}}{\sum^{c}_{{k=1}}\{e^{w_{k}}\}}\,,$ (18)
@@
${\mathcal{L}}^{\prime}_{r}=\sum_{t=1}^{c}\big{\{}\tilde{w}_{j}{\mathcal{L}}^{e}_{{\rm M}}(P_{{\bf Z}^{t}_{i}},P_{{\bf Z}^{t}_{i+1}})\big{\}}\,.$ (19)
@@
$w_{j}=w_{j}-\eta\nabla{\mathcal{L}}^{\prime}_{\rm total}\,,j=1,\cdots,c\,.$ (21)
@@ 3.5 Algorithm Implementation
The proposed framework is implemented in two main stages. First, we train the model on a sequence of tasks by minimizing a total loss function that consists of several components. The training procedure starts with the input dataset $D$, model parameters $\theta$, scalars $\alpha$, $\beta$, and $\gamma$, learning rate $\eta$, and the number of model layers $c$. Initially, the memory buffer $\mathcal{M}$ is empty, and the adaptive weight vector $\mathbf{w}$ is initialized with values sampled uniformly from a range of 0 to 1 for each feature layer. The training proceeds iteratively through each task, updating the total loss function at every step. During training, the model’s weights and the adaptive weights are optimized using gradients computed from the total loss function.
In each iteration of the algorithm, for each data point in the dataset, the model computes a total loss that includes three terms. The first term is the standard loss function $\mathcal{L}_{s}$, which evaluates the model’s prediction against the ground truth. The second term is a distance-based regularization, defined by the squared Euclidean distance between the current task’s feature representation and an exemplar from the memory buffer. The third term, which is computed only after the first task, incorporates the adaptive regularization term $\mathcal{L}^{\prime}_{r}$. This term quantifies the discrepancy between the feature layer outputs of the Teacher network (trained on the last task $T_{i-1}$) and the Student network (currently training on the current task $T_{i}$) by calculating the distance between the features $\mathcal{F_{t}}$ and $\mathcal{F_{s}}$ from corresponding layers of both networks. These distances are weighted by the normalized adaptive weights $\tilde{w}_{k}$, ensuring that each layer’s contribution to the regularization term is appropriately scaled.
The optimization of model parameters $\theta$ and adaptive weights $\mathbf{w}$ is performed using gradient descent, where the gradients of the total loss function are computed with respect to both the model parameters and the adaptive weights. The memory buffer $\mathcal{M}$ is updated using a reservoir sampling strategy [51], which stores the most relevant samples from the current task’s data. This procedure ensures that the memory buffer contains a balanced mix of data from both current and past tasks. The algorithm continues iteratively until all tasks have been processed, with the final model achieving continual learning across the entire sequence of tasks.
Algorithm 1 The training algorithm of the proposed MLFMM framework
1:  ${\bf Input:}$ dataset $D$, parameters $\theta$, scalars $\alpha$, $\beta$ and $\gamma$, learning rate $\eta$, number of model layers $c$
2:  ${\bf Init:}$$\mathcal{M}\leftarrow\{\}$, $\bf{w}\leftarrow Uniform(0,1,c)$
@@
4:     for $({\bf x},{\bf y})$ in $D^{s}_{i}$ do
5:        $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{s}(y,F_{\theta}({\bf x}))$
6:        ${\bf z}\leftarrow H_{\theta}({\bf x})$
7:        $(\bf{x^{\prime},z^{\prime}})\leftarrow sample(\mathcal{M})$
8:        $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{\rm total}+\alpha\cdot||{\bf z^{\prime}}-H_{\theta}({\bf x^{\prime}})||^{2}$
9:        $(\bf{x^{\prime},y^{\prime}})\leftarrow sample(\mathcal{M})$
10:        $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{\rm total}+\beta\cdot\mathcal{L}_{s}(y^{\prime},F_{\theta}({\bf x^{\prime}}))$
11:        if $i$ $>$then
12:           $\bf{x^{\prime}}\leftarrow sample(\mathcal{M})$
13:           $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{\rm total}+\gamma\cdot\mathcal{L}^{\prime}_{r}({\bf x^{\prime}})$
14:           $\theta\leftarrow\theta-\eta\nabla{\mathcal{L}}_{\rm total}$
15:           ${\bf w}\leftarrow{\bf w}-\eta\nabla{\mathcal{L}}_{\rm total}$
16:        else
17:           $\theta\leftarrow\theta-\eta\nabla{\mathcal{L}}_{\rm total}$
18:        end if
19:        $\mathcal{M}\leftarrow\textbf{reservoir}(\mathcal{M},({\bf x},{\bf y},{\bf z}))$
20:     end for
@@
Figure 1: The learning process of the first task, where the model learns without adaptive regularization.
@@
1:  ${\bf Input:}$ input data ${\bf x}$, adaptive weights $\bf w$
@@
7:  $\mathcal{F}_{t}\leftarrow{\bf x}$
8:  $\mathcal{F}_{s}\leftarrow{\bf x}$
9:  for $k=1$ to $c$ do
10:     ${\mathcal{F}}_{t},{\mathcal{F}}_{s}\leftarrow F^{\rm feature}_{\theta_{i-1},k}({\mathcal{F}}_{t}),F^{\rm feature}_{\theta_{i},k}({\mathcal{F}}_{s})$
11:     $\tilde{w}_{k}\leftarrow e^{w_{k}}/w_{sum}$
12:     $d\leftarrow d+\tilde{w}_{k}\cdot{\mathcal{L}}^{e}_{\rm M}({\mathcal{F}}_{t},{\mathcal{F}}_{s})$
13:  end for
14:  ${\bf Output:}$ $d$
As shown in Figure 1, the process for the first task involves the model learning from the initial task without applying the adaptive regularization term. During this phase, the model focuses solely on the current task’s loss function without considering past tasks. In subsequent tasks ($i>1$), as depicted in Figure 2, the model begins to incorporate the adaptive regularization term $\mathcal{L}_{r}$, which helps prevent forgetting while learning new tasks. This process ensures that the model retains knowledge from previous tasks by leveraging adaptive weights that regulate the influence of each layer in the network. Figure 3 illustrates how the $\mathcal{L}_{r}$ term is computed, showing the weighted distances between feature representations from the current and previous tasks, which are then incorporated into the overall loss function to maintain stability in learning.
Figure 2: The learning process of the subsequent tasks, where adaptive regularization is applied to the primary loss function.
Figure 3: Illustration of the computation of the adaptive regularization term $\mathcal{L}_{r}$, showing the weighted distances between feature representations across tasks.

### v4

@@ 3.1 Problem definition
$F_{\theta_{i},k}$ The $k$-th feature layer of $F_{\theta_{i}}$.
$G_{\theta_{i}}$ Linear classifier of $F_{\theta_{i}}$.
${\mathcal{Z}}^{k}$ Feature space of the outputs of $F_{\theta_{i},k}$.
$H_{\theta_{i}}$ Logits output: $H_{\theta_{i}}(\mathbf{x})=G_{\theta_{i}}(F_{\theta_{i},K}(\cdots F_{\theta_{i},1}(\mathbf{x})))$
$\mathcal{Z}^{c}$ Output space of classifier $G_{\theta_{i}}$ (logits space)
$\mathcal{F}({\theta_{i}},\cdot,k)$ The feature vector extracted by $F_{\theta_{i},k}$.
@@
$\mathcal{L}_{r}$ Multi-level regularization term using MMD: $\sum_{k=1}^{K}\tilde{w}_{k}\mathcal{L}^{e}_{\rm M}(P_{\mathbf{Z}^{k}_{i}},P_{\mathbf{Z}^{k}_{i+1}})$
$\mathcal{L}^{e}_{\rm M}(\cdot,\cdot)$ Unbiased MMD estimator between feature distributions (Eq. 13)
@@
$w_{k},\tilde{w}_{k}$ Adaptive weight for layer $k$ and its softmax-normalized version: $\tilde{w}_{k}=e^{w_{k}}/\sum_{j}e^{w_{j}}$
@@ 3.2 Multi-Level Feature Matching Mechanism
In the realm of continual learning, numerous studies have advocated for the utilization of knowledge distillation [54, 55, 39] methodologies to mitigate the phenomenon of network forgetting. The core principle of knowledge distillation involves maintaining the previous model as a teacher module and synchronizing the outputs between the teacher and the student module that is represented by the current active classifier. This alignment seeks to minimize substantial alterations in critical parameters of the current model (student) during the acquisition of new tasks [56]. Nevertheless, the majority of existing knowledge distillation techniques primarily focus on deriving a regularization term within the prediction space, often neglecting the semantic feature space. In this paper, we present an innovative approach termed the Multi-Level Feature Matching Mechanism (MLFMM), designed to regulate information flow within the feature space, which can effectively avert network forgetting.
@@
Let us formally define the function $F_{\theta_{i}}\colon{\mathcal{X}}\to{\mathcal{Y}}$ as a model parameterized by $\theta_{i}$, which comprises $K$ feature layers denoted as $\{F_{\theta_{i},1},F_{\theta_{i},2},\cdots,F_{\theta_{i},K}\}$, along with a linear classifier represented by $G_{\theta_{i}}\colon{\mathcal{Z}}\to{\mathcal{Y}}$. Here, the index $i$ signifies the model’s adaptation during the $i$-th task learning phase, and ${\mathcal{Z}}$ refers to the feature space corresponding to the output of the classifier. For a specified input ${\bf x}$, we can derive the feature vector from a particular feature layer through the following process :
@@
$\mathcal{F}({\theta_{i}},{\bf x},k)=\begin{cases}F_{\theta_{i},1}({\bf x})&\text{if }k=1\\ F_{\theta_{i},k}\left(F_{\theta_{i},1}({\bf x})\right)&\text{if }k=2\\ F_{\theta_{i},k}\left(F_{\theta_{i},k-1}\left(\cdots F_{\theta_{i},1}({\bf x})\right)\right)&\text{if }3\leq k\leq K\,,\end{cases}$ (2)
@@
where each feature layer $F_{\theta_{i},k}$ receives the output from the last layer $F_{\theta_{i},k-1}$ and produces the feature vector over the space ${\mathcal{Z}}^{k}$. The output of the classifier, or the logits, denoted as ${\bf z}\colon{\bf z}\in{\mathcal{Z}}$, is computed by applying the linear transformation $G_{\theta_{i}}$ to the final feature representation from the $K$-th layer:
${\bf z}=H_{\theta_{i}}(\mathbf{x})=G_{\theta_{i}}(F_{\theta_{i},K}(F_{\theta_{i},K}(\cdots F_{\theta_{i},1}({\bf x}))))\,,$ (3)
The whole prediction for the model $F_{\theta_{i}}$ can be expressed as
${\bf\hat{y}}=F_{\theta_{i}}({\bf x})\triangleq softmax({\bf z})\,,$ (4)
where ${\bf\hat{y}}$ is the prediction for the input ${\bf x}$ and $C$ is the total number of classes. The $softmax()$ is the normalized function. For logits ${\bf z}=\{z_{1},\cdots,z_{C}\}$, we have :
$z_{t}=\frac{e^{z_{t}}}{\sum_{j=1}^{C}e^{z_{j}}},t=1,\cdots.C,$ (5)
@@
In this paper, we adopt the cross-entropy loss function to update the parameter ${\theta_{i}}$, defined by :
${\mathcal{L}}_{s}({\bf y},{\bf\hat{y}})=\sum_{t=1}^{C}\big{\{}{\bf y}[t]\log({\bf\hat{y}}[t])\big{\}}\,,$ (6)
@@
To mitigate the issue of network forgetting, this study implements a memory buffer ${\mathcal{M}}_{i}$ to retain a limited number of historical examples. In particular, we propose utilizing Reservoir Sampling for the sample selection process, which offers computational efficiency. Nonetheless, the model’s performance on prior tasks diminishes when the memory capacity is constrained. To tackle this challenge, the proposed OWMMD regulates the updates to the model’s representations during the learning of new tasks. Specifically, upon the completion of the current task $T_{i}$, we preserve and freeze the parameters of all feature layers $\{F_{\theta_{i},1},\cdots,F_{\theta_{i},K}\}$ during the subsequent task learning $T_{i+1}$. A regularization term is incorporated to minimize the discrepancy in representations between the previous model ${F_{\theta_{i}}}$ and the current model ${F_{\theta_{i+1}}}$ at $T_{i+1}$ :
${\mathcal{F}}_{r}({\theta_{i}},{\theta_{i+1}},{\bf x})=\sum_{k=1}^{K}\big{\{}F_{d}({\mathcal{F}}({\theta_{i}},{\bf x},k),{\mathcal{F}}({\theta_{i+1}},{\bf x},k)\big{\}}\,,$ (7)
@@
where $F_{d}(\cdot,\cdot)$ is a distance measure function and $K$ is the total number of feature layers. In the following section, we introduce a probabilistic distance to implement Eq. (7).
@@ 3.3 The Maximum Mean Discrepancy based Regularization
${\bf Z}^{j}_{i}=\{\mathcal{F}({\theta_{i}},{\bf x}_{1},j),\cdots,\mathcal{F}({\theta_{i}},{\bf x}_{b},j)\}\,,$ (15)
Similar, we can form the feature vector ${\bf Z}^{j}_{i+1}$ using the current model $F_{\theta_{i+1}}$ on the data batch $\{{\bf x}_{1},\cdots,{\bf x}_{b}\}$, expressed as :
${\bf Z}^{j}_{i+1}=\{\mathcal{F}({\theta_{i+1}},{\bf x}_{1},j),\cdots,\mathcal{F}({\theta_{i+1}},{\bf x}_{b},j)\}\,.$ (16)
Let $P_{{\bf Z}^{j}_{i}}$ and $P_{{\bf Z}^{j}_{i+1}}$ denote the distribution of ${\bf Z}^{j}_{i}$ and ${\bf Z}^{j}_{i+1}$, respectively. Based on the MMD criterion and Eq. (7), we can implement the regularization function when seeing the data batch $\{{\bf x}_{1},\cdots,{\bf x}_{b}\}$ at the new task learning $T_{i+1}$ as :
${\mathcal{L}}_{r}=\sum_{t=1}^{K}\big{\{}{\mathcal{L}}^{e}_{{\rm M}}(P_{{\bf Z}^{t}_{i}},P_{{\bf Z}^{t}_{i+1}})\big{\}}\,.$ (17)
@@
where $P_{D^{s}_{i+1}}$ and $P_{{\mathcal{M}}_{i+1}}$ denote the distribution of $D^{s}_{i+1}$ and the memory buffer ${\mathcal{M}}_{i+1}$ at the $(i+1)$-th task learning, ${\bf z}$ is the logits of previous task and was saved in the memory buffer. This approach helps the model retain past knowledge while learning new tasks efficiently.
Since the proposed approach can be smoothly applied to the existing continual leanring models, we consider applying our approach to DER++. The final objective function, including a regularization term $\mathcal{L}_{r}$, is defined as :
@@ 3.4 Adaptive Regularization Optimization Term
To improve the performance of the proposed MMD-based regularization method [57, 61], it is essential to automatically allocate distinct weights to each regularization term. This strategy enables the assignment of varying degrees of significance to the features derived from different layers, thereby accurately reflecting their contributions to the overall learning framework. Specifically, we can establish an adaptive weight for each layer to assess its relevance to the entire learning process. To ensure that the adaptive weights across all layers are normalized and suitably balanced, we incorporate the softmax function.
Let ${\bf w}=\{w_{1},\cdots,w_{K}\}$ be a trainable adaptive weight vector, where $w_{j}$ determines the importance of the $j$-th feature layer. To avoid numerical overflow, we propose to normalize all weights using the following equation :
$\tilde{w}_{j}=\frac{e^{w_{j}}}{\sum^{K}_{{k=1}}e^{w_{k}}}\,,$ (20)
@@
${\mathcal{L}}^{\prime}_{r}=\sum_{t=1}^{K}\big{\{}\tilde{w}_{j}{\mathcal{L}}^{e}_{{\rm M}}(P_{{\bf Z}^{t}_{i}},P_{{\bf Z}^{t}_{i+1}})\big{\}}\,.$ (21)
@@
$w_{j}=w_{j}-\eta\nabla{\mathcal{L}}^{\prime}_{\rm total}\,,j=1,\cdots,K\,.$ (23)
@@ 3.5 Algorithm Implementation
The model training process of the proposed OWMMD framework follows a sequence of steps designed to incrementally learn from new tasks while mitigating catastrophic forgetting. The process is divided into distinct stages for each task, and the incorporation of adaptive regularization enhances the model’s ability to balance knowledge retention and learning of new tasks. The whole training algorithm of the proposed OWMMD framework is summarized into three steps :
Figure 1: The learning process of the proposed framework during the first task, where the model learns data samples without adaptive regularization. Images of Current Task Dataset and their logits are stored into the buffer via reservoir sampling for future task regularization. Different colored arrows indicate distinct data paths.
Figure 2: The learning process of the proposed framework during the subsequent tasks, where the adaptive regularization is applied to relieve network forgetting. The regularization term leverages layer-wise Maximum Mean Discrepancy (MMD) between the intermediate features of the teacher network (previous task model) and the student network (current model), guided by adaptive weights to prioritize critical layers. Current task data are processed to generate logits, which are then stored into the buffer via reservoir sampling, ensuring balanced retention of historical knowledge while adapting to new tasks. In the figure, Different colored arrows indicate distinct data paths.
Algorithm 1 The Training Algorithm of the Proposed OWMMD Framework
1:  ${\bf Input:}$ dataset $D^{s}$, parameters $\theta$, scalars $\alpha$, $\beta$ and $\gamma$, learning rate $\eta$, number of model layers $K$
2:  ${\bf Init:}$${\mathcal{M}}\leftarrow\{\}$, ${\bf w}\leftarrow Uniform(0,1,K)$
@@
4:     $\theta_{i}\leftarrow\theta$, ${\mathcal{M}_{i}}\leftarrow{\mathcal{M}}$
5:     for $({\bf x},{\bf y})$ in $D^{s}_{i}$ do
6:        $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{s}({\bf y},F_{\theta_{i}}({\bf x}))$
7:        ${\bf z}\leftarrow H_{\theta_{i}}({\bf x})$
8:        $({\bf x^{\prime}},{\bf z^{\prime}})\leftarrow sample({\mathcal{M}_{i}})$
9:        $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{\rm total}+\alpha\cdot||{\bf z^{\prime}}-H_{\theta_{i}}({\bf x^{\prime}})||^{2}$
10:        $({\bf x^{\prime}},{\bf y^{\prime}})\leftarrow sample({\mathcal{M}_{i}})$
11:        $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{\rm total}+\beta\cdot\mathcal{L}_{s}({\bf y^{\prime}},F_{\theta_{i}}({\bf x^{\prime}}))$
12:        if $i$ $>$then
13:           ${\bf x^{\prime}}\leftarrow sample({\mathcal{M}_{i}})$
14:           $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{\rm total}+\gamma\cdot{\mathcal{L}}^{\prime}_{r}({\bf x^{\prime}})$
15:           $\theta_{i}\leftarrow\theta_{i}-\eta\nabla{\mathcal{L}}_{\rm total}$
16:           ${\bf w}\leftarrow{\bf w}-\eta\nabla{\mathcal{L}}_{\rm total}$
17:        else
18:           $\theta_{i}\leftarrow\theta_{i}-\eta\nabla{\mathcal{L}}_{\rm total}$
19:        end if
20:        ${\mathcal{M}_{i}}\leftarrow reservoir({\mathcal{M}}_{i},({\bf x},{\bf y},{\bf z}))$
@@
22:     $\theta\leftarrow\theta_{i}$, ${\mathcal{M}}\leftarrow{\mathcal{M}_{i}}$
23:  end for
@@
1:  ${\bf Input:}$ adaptive weights $\bf w$
@@
7:  ${\bf x^{\prime}}\leftarrow sample({\mathcal{M}})$
8:  $\mathcal{F}_{t}\leftarrow{\bf x^{\prime}}$
9:  $\mathcal{F}_{s}\leftarrow{\bf x^{\prime}}$
10:  for $k=1$ to $K$ do
11:     ${\mathcal{F}}_{t},{\mathcal{F}}_{s}\leftarrow F_{\theta_{i-1},k}({\mathcal{F}}_{t}),F_{\theta_{i},k}({\mathcal{F}}_{s})$
12:     $\tilde{w}_{k}\leftarrow e^{w_{k}}/w_{sum}$
13:     $d\leftarrow d+\tilde{w}_{k}\cdot{\mathcal{L}}^{e}_{\rm M}({\mathcal{F}}_{t},{\mathcal{F}}_{s})$
14:  end for
15:  ${\bf Output:}$ $d$
Step 1: Initialization and Setup. The training process begins with initializing the dataset $D^{s}$, the model parameters $\theta$, and essential hyperparameters $\alpha$, $\beta$, and $\gamma$, which control the respective loss terms. A uniform weight vector $\mathbf{w}$ is also initialized for each layer of the model, and the memory buffer $\mathcal{M}$ is set as an empty set. These initializations provide the foundation for the learning process.
Step 2: Adaptive Regularization Optimization. The model optimizes the adaptive regularization term to ensure the stability of learned features across tasks, as showed in Figure 2. A key component of this step is the calculation of the distance between the teacher and student models at each layer. Specifically, during training, the teacher model ($\theta_{i-1}$) and the student model ($\theta_{i}$) are compared using the MMD-based regularization method, which calculates the distance between the features produced by the models at each layer. The detailed layer-wise MMD computation and its integration with adaptive regularization are described in Algorithm 1.
The adaptive weights play a critical role in this optimization. Each normalized weight, $\tilde{w}_{k}$, adjusts the importance of the corresponding feature layer, ensuring that the regularization process focuses on the layers that are most relevant to the current learning task. The softmax-normalized weight vector ensures that the contributions from all layers are appropriately balanced.
Furthermore, the distance between the teacher and student features, denoted as $\mathcal{L}^{e}_{\rm M}({\mathcal{F}}_{t},{\mathcal{F}}_{s})$, is computed at each layer to relive network forgetting. This mechanism ensures that the student model retains the essential knowledge from previous tasks while adapting to the new data. The final regularization loss integrates these distance terms across all layers, weighted by the adaptive weights.
Step 3: The Training Process. The overall training process is detailed in Algorithm 2. During the first task, as depicted in Figure 1, the model learns from the initial task without applying the adaptive regularization term. In this phase, the model focuses solely on the current task’s loss function. Notably, images from the current task dataset and their corresponding logits are stored into the buffer via reservoir sampling for future regularization. In the figure, different colored arrows indicate distinct data paths.
For subsequent tasks ($i>1$), as shown in Figure 2, the adaptive regularization term $\mathcal{L^{\prime}}_{r}$ is incorporated to mitigate network forgetting. In this stage, the regularization term leverages the layer-wise Maximum Mean Discrepancy (MMD) between the intermediate features of the teacher network (previous task model) and the student network (current model). This process is guided by adaptive weights to prioritize the critical layers. Additionally, current task data are processed to generate logits, which are then stored into the buffer via reservoir sampling, ensuring balanced retention of historical knowledge while adapting to new tasks.
The model continuously trains over all tasks in the sequence, updating both the adaptive weights $\mathbf{w}$ and the model parameters $\theta_{i}$ during each task learning phase. After each task, the model’s performance is evaluated and the memory buffer is updated with new data samples for future use, allowing the model to progressively generalize across tasks while preserving knowledge from previous ones.